OpenStudy (anonymous):

what is 5/6 +5 -4/8= i am lost with that

OpenStudy (anonymous):

yes it is

OpenStudy (sandra):

ok so first let's try reducing the individual parts, e.g. -4/8 = -1/2

OpenStudy (sandra):

so now you want to find the lowest common denominator for 5/6, 5, and -1/2

OpenStudy (sandra):

to me that looks like 6

OpenStudy (sandra):

so now you want to convert everything to sixths so you can add them up

OpenStudy (sandra):

so 5/6 is done

OpenStudy (sandra):

5 = 5*6/6 = 30/6

OpenStudy (sandra):

and -1/2 = -1/2 * 3/3 = -3/6

OpenStudy (sandra):

so now let's add it up: 5/6 +30/6 -3/6 = (5+30-3)/6 = 32/6

OpenStudy (sandra):

now you want to reduce that fraction, so again we need a common divider

OpenStudy (sandra):

looks like the only common divisor is 2

OpenStudy (sandra):

so 32/6 = 16/3

OpenStudy (sandra):

now if you want to , you can change that into a mixed fraction

OpenStudy (sandra):

so 16/3 = 15/3 + 1/3 = 5 and 1/3

OpenStudy (anonymous):

ok i am back

OpenStudy (sandra):

hey! wb

OpenStudy (sandra):

did that make sense?

OpenStudy (anonymous):

srry my cumputer is giving me problemes agine yes it did thanx

OpenStudy (sandra):

ok great =)

OpenStudy (sandra):

the basic point is, anything x 1 = the same thing

OpenStudy (sandra):

so you can always multiple by various forms of 1 to get a common denominator

OpenStudy (anonymous):

ok

OpenStudy (sandra):

so in the above, since I decided we needed a 6 on the bottom of all the fractions (common denominator), I multiplied 5 by 6/6 , and -1/2 by 3/3

OpenStudy (sandra):

to get 30/6 and -3/6

OpenStudy (sandra):

to reduce fractions, you're still essentially multiplying by 1

OpenStudy (sandra):

so to reduce 32/6, I divided both the top and bottom by 2

OpenStudy (anonymous):

how dothen do u flip it

OpenStudy (anonymous):

sorry then do u flip it

OpenStudy (sandra):

so how do you mean flip it

OpenStudy (sandra):

reducing fractions is still multiplying (or dividing if you will) by 1

OpenStudy (sandra):

just you're making the numbers smaller instead of bigger

OpenStudy (anonymous):

32 is on top then 6 is on the bottem you flip it to were the 6 is on top and the 32 is on the bottem

OpenStudy (sandra):

in reality, it's easiest to think "if I divide both the top and bottom of the fraction by the same thing, it will still be equivalent"

OpenStudy (anonymous):

oh ok that make senc

OpenStudy (anonymous):

i am not good at math

OpenStudy (sandra):

it's ok, nobody starts out good at math. takes lots of practice and not giving up on it

OpenStudy (sandra):

so let's do another example: reduce 48/12

OpenStudy (anonymous):

ok

OpenStudy (sandra):

your thought process should be: "what is the biggest number that evenly divides into both the top and the bottom"

OpenStudy (sandra):

and then do that

OpenStudy (sandra):

and it's ok if you don't know off the top of your head what the biggest number is

OpenStudy (sandra):

that also comes with practice

OpenStudy (sandra):

but you at least know in the above case, since bot the top and bottom are even, you can divide both of them by 2

OpenStudy (sandra):

so go ahead and try to reduce 48/12

OpenStudy (sandra):

I'll do a first step

OpenStudy (sandra):

48/12 = (48/2) / (12/2) = 24/6. I know that's true because I divided by 2/2, which is 1, and anyting divided by 1 = itself

OpenStudy (anonymous):

ok that is what i got to

OpenStudy (sandra):

ok, now the next question you should ask yourself is if there is another number you can divide 24/6 by to get a smaller fraction - by smaller I mean smaller numbers on the top and bottom

OpenStudy (sandra):

that process is called reducing

OpenStudy (anonymous):

i am geting so made at my comeputer right know

OpenStudy (anonymous):

if i log out agine i will get back on ok it is my computer

OpenStudy (sandra):

ok np

OpenStudy (anonymous):

we can divide 2 in to 24 and 6

OpenStudy (anonymous):

oh nvm not six

OpenStudy (anonymous):

and if it seems i am not i am i just in a livsseion i go to flvs

OpenStudy (anonymous):

but i will tell you be for i go tho ok

OpenStudy (sandra):

kk, well you were actually right - you *can* divide both 24 and 6 by 6 =)

OpenStudy (sandra):

so I would do that

OpenStudy (anonymous):

it would come ou to be 4/1

OpenStudy (anonymous):

i think not sure tho

OpenStudy (sandra):

you're right =) . and 4/1 is actually 4

OpenStudy (sandra):

good work =)

OpenStudy (sandra):

so how would you check to see if you're right?

OpenStudy (sandra):

an easy way to check is with a calculator

OpenStudy (sandra):

is 48/12 = 4/1 = 4?

OpenStudy (anonymous):

ok thanx you vary much

OpenStudy (sandra):

no problem!

OpenStudy (sandra):

so just remember - you can always multiply the top and the bottom by the same thing, to get an equivalent fraction. that's useful for getting a common denominator to make adding/subtracting fractions easier

OpenStudy (sandra):

and that's because you're *Actually* just multiplying by 1

OpenStudy (anonymous):

ok the way my techer did it was confusing

OpenStudy (sandra):

and anything times 1 is the same exact thing, just written in a different way

OpenStudy (sandra):

now when you're "reducing" fractions, all you're trying to do is get the top and bottom numbers as small as possible

OpenStudy (sandra):

for that, you can always divide both the top and bottom by the same thing, to get an equivalent fraction

OpenStudy (anonymous):

know with that rule does it count with 0 to

OpenStudy (sandra):

since in reality, you're just dividing tat number by 1

OpenStudy (sandra):

and anything divided by 1 is the same thing, just written in a different way

OpenStudy (sandra):

well, you can't divide by 0

OpenStudy (sandra):

or else the world explodes =)

OpenStudy (anonymous):

ok @};-

OpenStudy (anonymous):

is open study only for math

OpenStudy (sandra):

that I know of

OpenStudy (sandra):

well, I think there's another UI they have

OpenStudy (sandra):

but that's over at openstudy.com, and I don't think the logins are shared

OpenStudy (sandra):

I like this one more

OpenStudy (anonymous):

ok i think it should make it for all subjegs but i guess it dos not relly matter tho

OpenStudy (sandra):

what else are you taking?

OpenStudy (anonymous):

i am taking teck math langue arts chess and on the side i am lourning chines japens spanish french

OpenStudy (anonymous):

i mean tech for cumputers

OpenStudy (anonymous):

nehow

OpenStudy (anonymous):

means hello in chines

OpenStudy (anonymous):

2 to 4 i have a chess compition

OpenStudy (sandra):

oh awesome =)

OpenStudy (sandra):

that's a lot!

OpenStudy (anonymous):

yea but in reality i should lour alot more cuz i mean i am not big on politics but ii do know alot about how most school seystem r laking on some eguoction

OpenStudy (anonymous):

my friend is in 8th he deos not know anthing but it is also the pareints to not all on the schools

OpenStudy (anonymous):

the schools can only teach so much in won day and they can not stay on won topec for long

OpenStudy (anonymous):

i am not going to lie i use to be won of thos kide who were like when am i ever going to use any of this stuff but in reality u use it all the time

OpenStudy (anonymous):

even if you relis it or not